the perimeter of a rectangle is 36 yd. The width is 18 yd less than twice the length. find the length and width of the rectangle
Answer by mandi08(1)

Perimeter = 2W + 2L
36 = 2W + 2L
We will use L as our variable.
L is the length; the width is 18 less than twice the length, therefore (2L - 18).
Plug this in to the perimeter equation and you get:
36 = 2(2L-18) + 2(L)
36 = 4L - 36 + 2L
72 = 4L + 2L
72 = 6L
L=12
Now we know the length, so we need to figure out the width. The equation to find the width is:
W = 2L - 18
W = 2(12) - 18
W = 24 - 18
W = 6